### v2

# Design2GarmentCode: Turning Design Concepts to Tangible Garments Through Program Synthesis

Abstract
Sewing patterns, the essential blueprints for fabric cutting and tailoring, act as a crucial bridge between design concepts and producible garments. However, existing uni-modal sewing pattern generation models struggle to effectively encode complex design concepts with a multi-modal nature and correlate them with vectorized sewing patterns that possess precise geometric structures and intricate sewing relations. In this work, we propose a novel sewing pattern generation approach Design2GarmentCode based on Large Multimodal Models (LMMs), to generate parametric pattern-making programs from multi-modal design concepts. LMM offers an intuitive interface for interpreting diverse design inputs, while pattern-making programs could serve as well-structured and semantically meaningful representations of sewing patterns, and act as a robust bridge connecting the cross-domain pattern-making knowledge embedded in LMMs with vectorized sewing patterns. Experimental results demonstrate that our method can flexibly handle various complex design expressions such as images, textual descriptions, designer sketches, or their combinations, and convert them into size-precise sewing patterns with correct stitches. Compared to previous methods, our approach significantly enhances training efficiency, generation quality, and authoring flexibility. Our code and data will be publicly available.
Figure 1: Traditional sewing pattern generation approaches (top) use uni-modal models trained on synthetic datasets generated by parametric pattern-making programs (red arrow) to convert text or image prompts into vector-quantized patterns. These methods are resource-intensive and often yield oversimplified patterns with stitching errors. Our approach (bottom) utilizes large pre-trained LMMs to directly translate design concepts into parametric programs and configuration files (blue arrow), enabling dedicated, structurally correct pattern generation from multi-modal design inputs within a unified framework.
1 Introduction
While generative AI has significantly propelled creativity in fashion design, turning those design ideas into wearable realities remains a formidable challenge. Sewing patterns are the key components to bridge the gap between abstract design ideas and wearable realities. They are foundational blueprints that dictate the precise shapes and dimensions of fabric pieces, essential for assembling garments in both the physical and virtual fashion realms.
Traditionally, sewing patterns are drafted manually by professional pattern-makers with years of practice, making the process inefficient, error-prone, and unable to meet the growing demands for refinement and personalization in the fashion market. To this end, parametric pattern-making researches [8, 28, 30] and industrial solutions [4, 2, 1, 3] have emerged. These methods formalize the pattern-making process as geometric functions governed by parameters such as body measurements and design features, thereby accelerating the process by enabling pattern makers to generate sewing patterns through parameter adjustments instead of starting from scratch. However, creating these function templates is still complex, requiring not only advanced pattern-making skills but also geometric intuition, mathematical modeling knowledge, and coding abilities to translate pattern-making expertise into CAD programs. These technical barriers significantly restrict the widespread adoption of parametric pattern-making solutions.
Figure 2: (a) Despite prompts specifying diverse neckline types, DressCode [21] consistently produces only V-neck designs, indicating limited generation diversity. (b) SewFormer [39] often generates sewing patterns with self-intersecting panels, compromising pattern validity. (c) Stitching errors are also prevalent in Sewformer [39], as shown here where a pant side seam is mistakenly stitched to a shirt shoulder seam, resulting in draping failure.
Recently, several learning-based approaches for sewing pattern generation have been introduced. For instance, NeuralTailor [29] focuses on extracting sewing patterns from unstructured point clouds, DressCode [21] targets text-to-sewing pattern generation, and SewFormer [39] is designed for image-based sewing pattern generation. However, they are generally trained on paired design-sewing pattern data, necessitating large datasets to effectively capture the multi-modal nature of design concepts. Furthermore, sewing patterns require centimeter-level precision to ensure proper garment fit, which presents a significant challenge for neural networks that only provide statistical approximations of the true function based on their training data  [22, 14, 46]. As a result, these methods frequently generate oversimplified patterns with flawed geometry or stitches, potentially leading to draping failures (Figure 2).
In this paper, we present Design2GarmentCode, an innovative approach that leverages the generalization capabilities of vision-language foundation models to achieve multi-modal sewing pattern generation with minimal computational and data requirements. Unlike previous methods that directly synthesize vector-quantized patterns, Design2GarmentCode employs LMMs to learn the syntax of parametric pattern-making programs, translating design concepts into parameters and programs that can be executed to produce precise and structurally accurate sewing patterns. Design2GarmentCode combines a pre-trained Large Multimodal Model (LMM) as a design interpreter with a finetuned Large Language Model (LLM) as a program synthesizer. Specifically, the LLM is finetuned on code snippets from GarmentCode [30], a domain-specific language for constructing parametric sewing patterns. At runtime, the design interpreter extracts both topological and geometrical information from the design input by responding to a series of questions from the program synthesizer, which then generates garment programs and design configurations following GarmentCode syntax. Our method offers the following major contributions:
  • We introduce a novel modality-agnostic framework with an intuitive, intelligent interface capable of processing user design intentions across multiple modalities simultaneously by integrating pre-trained LMMs.
  • We present the first sewing pattern generation approach grounded in program synthesis, delivering fully interpretable, geometrically precise, and structurally accurate patterns through a more compact, semantically clear, and LLM-friendly representation.
  • Our framework benefits real-world production by enabling flexible pattern authoring through natural language or physical feedback, allowing precise customization and efficient creation of novel garment components, represented as parametric pattern-making programs.
  • Our approach requires only minimal fine-tuning of a pre-trained LLM and the training of a lightweight, text-conditioned transformer decoder, making it more efficient than existing vector-quantized sewing pattern generation models trained from scratch while offering superior generation quality and authoring flexibility.
2 Related Work
2.1 Garment Modeling with Sewing Patterns
Garment modeling and generation can be broadly classified into two categories: direct 3D garment generation (meshes) and sewing pattern generation, which are later draped onto human bodies via cloth simulation [38, 55, 52] or learning-based techniques [36, 34]. Direct 3D garment generation often relies on differentiable garment representations like unsigned distance fields [66, 64], shells [41], or Gaussian splatting [47]. However, it presents challenges in terms of both geometric accuracy and editability. On one hand, capturing fine garment details like folds and wrinkles necessitates extremely high-resolution 3D representations. On the other hand, editing these generated garments requires a well-defined UV space, and flattening the 3D mesh into developable meshes demands careful consideration from both geometric and statistical perspectives [45, 7].
Sewing pattern generation, by contrast, has been approached through both learning-based and procedural modeling methods. Learning approaches utilize vector-quantized representations of sewing patterns, mapping from unstructured 3D point clouds [7, 29], images [25, 12, 39, 63], or textual descriptions [21] to structured patterns. However, they are highly dependent on the quality and diversity of the training data and often struggle to generalize designs beyond the training domain. Furthermore, generating high-quality 3D garment data requires substantial domain knowledge and the involvement of skilled professionals.
Procedural modeling is an alternative that relies on predefined rules and parameters to generate garment patterns. For instance, GarmentCode [30], a DSL for parametric pattern making, enables precise control over garment design and customization. The GarmentCodeData [31], built on GarmentCode, further illustrates the potential of procedural methods to generate a diverse range of made-to-measure garments, with adaptability to different body shapes. While procedural modeling provides greater control and precision, it typically requires specialized expertise and is less flexible when dealing with novel or unconventional designs.
2.2 LLMs for Program Synthesis
Recent advancements in program synthesis and code generation using large language models (LLMs) have laid essential groundwork for systems like Design2GarmentCode, which generate structured garment code from multi-modal design inputs. Earlier researches like Codex [11] and AlphaCode [35] demonstrated the effectiveness of LLMs in generating complex, task-specific code with high syntax accuracy, showcasing potential in scenarios requiring precise parametric coding. These models [42, 58] highlight how LLMs, when sufficiently trained, can transform natural language inputs into executable code, a capability directly relevant to generating garment codes that follow complex pattern-making syntax.
Additionally, researchers [54, 18, 9, 19, 62] have been exploring explore multi-modal models that integrate visual aids, such as flowcharts and UML diagrams, into LLM training to enhance models’ comprehension of complex structures and flow. These models particularly emphasize the need for semantic understanding and adaptability, which are critical in working with domain-specific languages (DSLs) like GarmentCode.
2.3 Neurosymbolic Models
Procedural/symbolic models and learned/neural models have complementary strengths and weaknesses. Neurosymbolic models [46] tend to combine the strengths of both paradigms and propose to generate visual data using symbolic programs augmented with AI/ML techniques. The neurosymbolic pipeline typically includes task specification, program synthesis using a DSL, program execution, and optional neural post-processing for refining results. It has been successfully applied across several areas of computer graphics. In 2D shape modeling, they are used in layout generation [53, 44], engineering sketch creation [15, 43, 48], and vector graphics synthesis by constructing programs that represent geometric shapes and their spatial relations [10, 16]. In 3D shape modeling, they facilitate inferring shape programs from existing 3D models [50, 32, 26, 27] or generating entirely new 3D shapes by training generative models on shape programs [56, 59, 60] or generate generate node graphs that define complex textures and materials [49, 20, 24] following the procedural modeling paradigm. Additionally, neuro-symbolic methods have been employed in human motion prediction [17, 37], reasoning [57, 33, 65] and generation [13, 40, 61], which leveraging visual-language foundation models to extract symbolic representations from visual data, facilitating complex activity reasoning by combining visual cues with symbolic logic [57].
Our approach aligns with the neurosymbolic paradigm by instruction-tuning LLMs to generate GarmentCode from various forms of design concepts. Similar to neurosymbolic models used in 2D/3D shape modeling and procedural texture generation, our method synthesizes structured, executable programs that define garment components, their relations, and parameters. This allows for precise, customizable generation of sewing patterns given multi-modal design inputs while using the symbolic power of GarmentCode to ensure geometric and structural accuracy.
Figure 3: Overview of Dress2GarmentCode. (1) Program Learning: we finetune the DSL Generation Agent (DSL-GA) using GarmentCode example programs, teaching it the GarmentCode grammar and the semantics of each design parameter. (2) Prompt Synthesis: the DSL-GA generates prompts for the Multi-Modal Understanding Agent (MMUA) to interpret and extract relevant design features from the input (3). (4) Program Synthesis: based on the MMUA’s responses, the DSL-GA synthesizes GarmentCode-compliant design configurations and garment programs, which are then executed by the GarmentCode engine to produce sewing patterns and simulated garments (5). To enhance robustness, we incorporate two validation loops: during program synthesis, we employ rule-based validations (7) to ensure the MMUA’s outputs are sufficient for generating complete and valid garment programs and design parameters; after the initial generation, the MMUA compares the generated design with the input and suggests modifications to minimize discrepancies.
3 Method
Our goal is to develop a generative model that transforms multi-modal design concepts into precise sewing patterns. This requires understanding diverse inputs and producing patterns with high geometric precision and intricate structures. These requirements present a challenge for conventional models, which require extensive training data and struggle with output precision due to their probabilistic nature. We propose Design2GarmentCode, a system leveraging LMMs to generate parametric pattern-making programs, or specifically GarmentCode [30]. Design2GarmentCode reduces the need for large datasets utilizing the pre-embedded pattern-making knowledge in LMMs while ensuring output precision with parametric program synthesis. In the following, we first provide an overview of parametric pattern-making programs and GarmentCode syntax, and then describe the detailed design of Design2GarmentCode.
3.1 Parametric Sewing Patterns
Parametric sewing patterns
are formally represented as symbolic programs that generate sewing patterns (i.e., 2D CAD sketches) based on body measurements and design configurations. These symbolic programs enhance the efficiency of the pattern-making process by allowing users to draft or modify sewing patterns through semantically meaningful parameters. Mathematically, we can represent a sewing pattern $\mathcal{S}$ as:
$\mathcal{S}=\langle\mathcal{F},\mathcal{D},B\rangle=\cup_{f_{i}\in\mathcal{F},d_{i}\in\mathcal{D}}f_{i}(d_{i},B),$ (1)
where $\mathcal{F}$ is the set of symbolic programs, $\mathcal{D}$ represents design configurations, and $B$ represents body measurements. Each symbolic function $f_{i}\in\mathcal{F}$ is essentially a series of rule-based 2D draw-calls controlled by its unique set of design configurations $d_{i}\in\mathcal{D}$ and the body measurements $B$.
GarmentCode
is a domain-specific language (DSL) designed to generate parametric sewing patterns by encapsulating those symbolic programs in a hierarchical, component-oriented manner. In GarmentCode, each symbolic program $f_{i}$ uses parametric curves to define a garment component, such as sleeves, bodices, or collars. The smallest component is a single panel, and multiple components can be combined through interface functions to create a larger component. In GarmentCode, a complete sewing pattern is specified by topological parameters $\mathcal{D}_{T}$ (which define the presence and quantity of garment components) and geometrical parameters $\mathcal{D}_{G}$, which determine the dimensions of each component when combined with body measurements $B$. As this work primarily focuses on design variations, we use a standard body model throughout all experiments to ensure consistency.
3.2 The Design2GarmentCode System
As illustrated in Figure 3, Design2GarmentCode has three components: DSL Generation Agent (DSL-GA), a finetuned LLM responsible for (1) program learning, (2) prompt synthesis, and (4) program synthesis; Multi-modal Understanding Agent (MMUA), a pre-trained LLM that manages design understanding (3) and design comparison (6); and (5) GarmentCode, which executes the synthesized programs to generate sewing patterns and 3D garments.
The system workflow begins with Program Learning (Sec. 3.2.1), where DSL-GA is finetuned to understand the syntax and semantic meanings of GarmentCode parameters. In Prompt Synthesis, DSL-GA creates prompts for MMUA to identify essential design features. These features are then provided to DSL-GA for Program Synthesis, where garment programs and design configurations are generated through rule-based parameter validation (Figure 3 (7)) and a learned projector (Sec. 3.2.2). The GarmentCode Execution Engine then produces sewing patterns and draped garment models. Finally, a Validation stage compares the generated garment with the original design, allowing MMUA to provide specific correction instructions to DSL-GA for iterative refinement, such as “make the sleeve longer”.
3.2.1 Program Learning
During experiments, we found that pre-trained LLMs have some foundational knowledge of pattern drafting. For example, when prompted with “How to draft a basic upper body bodice?”, LLMs can produce drafting instructions that align with conventional practices. We use a pre-trained LLM to initialize DSL-GA, however, due to GarmentCode’s customized object notations and function logistics, directly prompting DSL-GA to generate GarmentCode programs poses significant challenges [9].
To address these challenges, we propose to align DSL-GA’s embedded pattern-making knowledge with the specific syntax and semantics of GarmentCode via LoRA [23] based on fine-tuning. We start by providing the DSL-GA (denoted as $\Gamma$) with existing GarmentCode programs $\mathcal{F}$, and instructing it to comment on the functions with detailed pattern-drafting instructions. After manually validating the comments, we get a dataset $D$ paring natural language instructions with GarmentCode implementations:
$D=\bigg{\{}\big{(}\Gamma_{cmt}(f_{i}),f_{i}\big{)}\,\ |\,\,f_{i}\in\mathcal{F}\bigg{\}},$ (2)
where $\Gamma_{cmt}$ is the instructed DSL-GA for code commenting, $f_{i}$ is . Similar to [9], we finetune DSL-GA ($\Gamma$) on the dataset $D$ with LoRA [23], aiming for $\Gamma_{ft}(\Gamma_{cmt}(f_{i}))\rightarrow f_{i}$, where $\Gamma_{ft}$ is the finetuned DSL-GA.
After fine-tuning, the fine-tuned DSL-GA $\Gamma_{ft}$ gains an understanding of the code structure and parameter semantics in GarmentCode (Figure 9). Therefore, we provide the design configuration $\mathcal{D}$ to $\Gamma_{ft}$, prompting it to analyze the semantic meaning of each parameter and generate structured queries. These queries are designed to guide the MMUA in extracting relevant design features from the multi-modal input, enabling $\Gamma_{ft}$ to generate a comprehensive set of design parameters. The generated prompt $P$ typically starts with analysis instructions, followed by multiple-choice or numerical estimation questions regarding each design parameter $d_{i}$. Formally, we have
$P=\Gamma_{ft}(\mathcal{D})=\cup_{d_{i}\in\mathcal{D}}\Gamma_{ft}(d_{i})=\cup q_{i},$ (3)
where $q_{i}=\Gamma_{ft}(d_{i})$ represents the generated question regarding the $i$-th design parameter $d_{i}$.
3.2.2 Program Synthesis
Initial results showed that MMUA performed significantly better on multiple-choice questions compared to numerical estimation questions. To improve accuracy, we replaced all numerical estimation questions in the initial prompt $P$ with equivalent multiple-choice questions with descriptive options such as “full length”, “half length”, or “three-quarter length”. We append a lightweight projector $\Psi$ after the finetuned DSL-GA $\Gamma_{ft}$ to transform these descriptive answers $\tau_{i}$ regarding the design input $x$ into precise geometrical parameters $d_{i}\in\mathcal{D}$ adhering to GarmentCode [30]:
$\Psi:\Gamma_{ft}(\cup\tau_{i})\rightarrow\mathcal{D},\,\text{ where }\tau_{i}=\textbf{MMUA}(q_{i},x).$ (4)
Inspired by DressCode [21], we implement the projector $\Psi$ as text-conditioned decoder-only transformer, where we design a type-based quantization function $\mathbf{Q}$ to convert the parameter list $\mathcal{D}$ into a token sequence $\mathcal{T}=\{t_{1},...,t_{N}\}$, where $N=|\mathcal{D}|$ denote the total number of design parameters. The quantization function $\mathbf{Q}$ operates as follows:
$t_{i}=\mathbf{Q}(d_{i})=\begin{cases}0/1,&\text{if }d_{i}\text{ is a boolean variable},\\ d_{i},&\text{if }d_{i}\text{ is an integer},\\ \lambda\cdot\textbf{Norm}(d_{i}),&\text{if }d_{i}\text{ is a floating number},\\ \textbf{Index}(d_{i},L),&\text{if }d_{i}\text{ is a selective variable}.\\ \end{cases}$ (5)
where $\lambda$ is a scaling factor indicating numerical precision. We use $\lambda=100$ to maintain centimeter-level precision.
As in Eq. 4, we use the finetuned DSL-GA $\Gamma_{ft}$ to encode the answers $\tau_{i}$ from MMUA and construct the condition input for $\Psi$ (we use an MLP to match the embedding dimension between $3,072$ in $\Gamma_{ft}$ and $128$ in $\Psi$). Notably, our token sequence length is fixed to the number of design parameters $|\mathcal{D}|=122$, regardless of the complexity of the pattern. This fixed-length representation is at least 10×10$\times$10 × compact than DressCode [21], whose sequence length is $1,500$ and scales with pattern complexity (see Sec. 4.1 for implementation details).
4 Experiments
4.1 Implementation Details
We use GPT-4V [6] for MMUA, and an instruction tuned version of Llama-3.2-3B for DSL-GA ($\Gamma$). The following sections contains the detailed explanation for the finetuned DSL-GA, and training details for the Projector $\Psi$.
4.1.1 Finetuning DSL-GA $\Gamma_{ft}$
To optimize the trade-off between computational cost and generation quality, we implemented the DSL generation agent (DSL-GA) using the Llama-3.2-3B-Instruct[51] model, fine-tuned over two epochs with LoRA (rank 16) and a learning rate of $5\times 10^{-4}$. All code generation experiments were conducted on a single NVIDIA GTX 4090. For multi-modal understanding tasks, GPT-4V was employed as the designated agent.
4.1.2 Training The Projector $\Psi$
The projector $\Psi$ is trained on the GarmentCodeData [31] dataset, which comprises approximately 115,000 garment samples draped on a standard A-pose body. We generate initial design descriptions for each sample using GPT-4V or rule-based inverse mapping from the ground truth design parameters for the sample, for example {minted}python if design.shirt.length.v ¿ 1.0: return ’shirt__length__long’ The token sequence length is fixed at $122$, which is equal to the number of design parameters in GarmentCode. The projection MLP and Transformer decoder are designed with feature dimensions of $128$. The MLP consists of $4$ intermediate layers, while the Transformer decoder includes $8$ layers. Training is conducted using the Adam optimizer with a learning rate of $5\times 10^{-4}$, a batch size of $16$, and completed on a single NVIDIA GTX 4090 within $10$ hours.
Notably, although we adopt a decoder-only Transformer architecture similar to DressCode, our innovative approach of quantifying sewing patterns through design parameters proves to be significantly more efficient and scalable. Specifically, with DressCode’s quantization scheme, the token sequence length is calculated as:
$L_{seq}=N_{p}\times(N_{e}\times L_{e}+\|R\|+\|T\|+N_{e}\times\|S\|)+2$
where $N_{p}$, $N_{e}$ denotes the maximum number of panels and edges respectively. $\|R\|=4$ is the length of rotation quaternions, and $\|T\|=3$ is the length of 3D translation vector. $\|S\|=4$ represents the per-edge stitching parameters containing a stitch tag and its existence indicator. $L_{e}$ represent the length of quantified edge vectors, which might be $6$ for cubic bezier curves and $4$ for quadratic bezier curves. Using GarmentCodeData as an example, to fully cover GarmentCode’s modeling space, the required sequence length under DressCode’s method would be $13,951$, with $N_{p}=N_{e}=37,L_{e}=6$, which will cost $\approx 1.5h$ to generate a single sewing pattern using DressCode, while our token sequence length is fixed at $122$.
Method   Text Guided Generation   Image Guided Generation
DressCode [21] Ours Sewformer [39] Ours
Quality SSR 84% 100% 65.33% 94%
Agreement 7.17% 79.83% 3.33% 88.67%
Aesthetic 9.50% 68.17% 5.33% 77%
Diversity # Panels $5.11_{\pm 2.76}$ $6.92_{\pm 9.63}$ $10.11_{\pm\textbf{19.53}}$ $\textbf{11.02}_{\pm 17.48}$
# Edges $5.48_{\pm 2.56}$ $6.84_{\pm 11.42}$ $5.79_{\pm 2.91}$ $6.24_{\pm 8.4}$
# Stitches $10.06_{\pm 10.51}$ $18.66_{\pm 74.58}$ $15.81_{\pm 34.97}$ $27.9_{\pm 96.61}$
Table 1: Quantitative comparisons between our method and SOTA sewing pattern generation methods in terms of generation quality and diversity. SSR is Simulation Success Rate, Agreement quantifies alignment with design prompts, and Aesthetic evaluates the visual appeal of the generated patterns. #Panels, #Stitches, and #Edges represent the average and variance (subscript) number of panels/stitches per pattern and edges per panel, respectively.
4.2 Quantitative Evaluation
We evaluate our proposed method against state-of-the-art sewing pattern generation approaches (DressCode [21] for text-guided and Sewformer [39] for image-guided generation) on Generation Quality and Generation Diversity.
Generation Quality is evaluated through three metrics: Simulation Success Rate (SSR), Agreement Score, and Aesthetic Score. The Simulation Success Rate (SSR) is calculated as the ratio of successfully simulated garments to the total number of generated sewing patterns, measuring the structural feasibility of the patterns. We prepared a dataset comprising 150 text prompts and 150 test images. For each sample, we generated sewing patterns using both our method and baseline methods, and simulated the patterns using GarmentCode’s simulation engine [30, 31] to compute the success rate. The Agreement and Aesthetic Scores were derived from a user study involving 30 professional pattern-makers. Each participant is asked to review 50 text and 50 image test samples. For each sample, we present the participants’ sewing patterns and simulated garments generated by our method and the baseline models, and ask them to rate each based on two criteria:
  • Agreement: the degree to which the generated pattern matched the design prompt.
  • Aesthetic Quality: the visual appeal and structural coherence of the generated pattern.
For each criterion, participants could express a preference for either our method or the baseline or indicate that both methods were “comparable”. We then calculated the Agreement and Aesthetic Scores as the percentage of times each option was chosen over the total number of tested samples.
Table 1 presents the results, showing that our method surpasses existing approaches in both SSR and user-evaluated Agreement and Aesthetic scores. For text-guided generation, our model achieves a perfect 100% SSR, notably higher than DressCode’s 84%. Additionally, our Agreement score of 79.83% and Aesthetic score of 68.17% far exceed DressCode’s respective scores of 7.17% and 9.5%. In image-guided generation, our method attains a 94% SSR, with an Agreement score of 88.67% and an Aesthetic score of 77%, significantly outperforming Sewformer. These enhancements highlight our model’s ability to generate sewing patterns that are both structurally precise and visually aligned with the design prompt.
Generation Diversity
is evaluated by analyzing the average number of panels (# Panels), edges (# Edges), and stitches (# Stitches) in the generated patterns. For text-guided generation, our method yields more intricate designs, with an average of 6.92 panels, 6.84 edges, and 18.66 stitches per pattern, compared to DressCode’s simpler outputs of 5.11 panels, 5.48 edges, and 10.06 stitches. In image-guided generation, our approach also demonstrates superior diversity, producing an average of 11.02 panels, 6.24 edges, and 27.9 stitches per pattern, compared to Sewformer’s averages of 10.11 panels, 5.79 edges, and 15.81 stitches. These results emphasize our model’s ability to capture and replicate subtle design variations, highlighting its robustness and adaptability across different design inputs.
Figure 4: Quality Comparison on Text-Guided Sewing Pattern Generation. For each design, we present the generated pattern using our method (left) alongside DressCode [21] (right), including front and back renderings of the draped garment. We highlight design elements accurately captured by our method but missed by DressCode [21] use red color in the input prompt.
Figure 5: Quality Comparison on Image-Guided Sewing Pattern Generation. We compare our method with Sewformer [39] on Internet-collected fashion photographs (left), and AI-generated design images without human models (right). The results indicate that our method successfully captures design details from diverse styles, producing sewing patterns that accurately reflect neckline (a, d), cuffs (a, e, g), darts (c, d), and asymmetry (f). In contrast, Sewformer’s results exhibit several issues, including incorrect necklines (a, d), missing components (b, g), misplaced or imaginary stitches (d, e), and extraneous pattern pieces (h). Additionally, since Sewformer’s pattern generation does not account for body shape, garments like skirts and pants frequently appear oversized around the waist, causing them to sag when draped.
Figure 6: Examples of sketch-based sewing pattern generation. Our method was able to generate high-quality sewing patterns from design sketches under various styles and could integrate seamlessly with industrial fashion design software for (a) pattern editing, i.e. sleeve panels in red boxes are merged from separate front/back sleeve panels; and (b) avatar posture and fabric material editing.
Figure 7: Sewing Pattern Authoring with instructions. Starting from an original design, the system follows user instructions to adjust specific pattern elements. In Edited Design 1, the pants are modified to a skirt based on the command “CHANGE THE PANT TO SKIRT”. In Edited Design 2, the sleeves are shortened as requested. Finally, in Edited Design 3, the shirt is made sleeveless in response to the instructions. Note that, each modification accurately applies only to the specified parts, leaving the rest of the design unchanged.
4.3 Multi-modal Generation Results
Our proposed method demonstrates superior performance across various sewing pattern generation tasks, including text-guided, image-guided, and sketch-based generation.
In text-guided sewing pattern generation (Figure 4), our method accurately captures design details specified in prompts, such as neckline types (e.g., crew neck (a), boat neck (b), turtleneck (c)) and complex structural features like asymmetry (f) and layered skirts (d). In comparison, the baseline model DressCode struggles with limited pattern diversity, often defaulting to simpler shapes like V-neck designs. Additionally, for design descriptions out of its training domain, DressCode frequently generates patterns with incorrect stitching, leading to poor draping results (Figure 4 (e)). Our method could provide structurally sound and visually accurate patterns under a large design variety, showcasing its capability to handle diverse design requests with high fidelity.
For image-guided sewing pattern generation (Figure 5), our model effectively translates detailed visual cues from input images into corresponding sewing patterns. Compared with Sewformer, which often fails to model-specific design elements like cuffs, hoods, and asymmetric features, our approach accurately reproduces these details. Sewformer’s results frequently exhibit structural flaws, such as missing or misaligned pattern pieces and extraneous components, resulting in unrealistic garment draping. In contrast, our method maintains structural integrity and captures complex design features, producing patterns that closely align with the source images.
In sketch-based sewing pattern generation (Figure 6), our system seamlessly converts both technical sketches (left) and artistic drawings [^1] The drawing is borrowed from the artwork of TWELVEYIN. (right) into high-quality sewing patterns. We also demonstrated that the generated sewing patterns could seamlessly integrate into industrial fashion design software [^2] We use Style3D Studio [5] for pattern and appearance authoring.. For example, highlighted sleeve panels in Figure 6 (a) are merged from separate front and back sleeve pieces, while Figure 6 (b) demonstrates avatar posture and fabric material editing.
5 Application
In this section, we explore practical applications enabled by our system that extend beyond basic pattern generation, providing designers with versatile tools for design refinement, integration with physical simulation, and the creation of new garment components.
Instruction-Based Editing. Our system allows designers to adjust generated sewing patterns through simple, instruction-based edits, utilizing the same refinement process as in our collaborative framework. As illustrated in Figure 7, starting from an original design, the system responds to natural language commands from the user to adjust the sewing pattern. At each step, the modified areas are highlighted in red boxes. From the figure, it is evident that our system can accurately update only the specified parts of the pattern according to the user’s instructions while leaving all other parts of the design unchanged.
Figure 8: Sewing pattern adjustment based on body pressure measurement. Red regions indicate areas of tight fabric with high body pressure, while blue regions represent looser areas.
Figure 9: The code for a layered-skirt component generated by our DSL-GA and 3D garment under various design parameters.
Physics-Based Editing. Our system’s generated sewing patterns integrate seamlessly with professional cloth simulation software, allowing adjustments based on fitness measurements derived from physical simulations. In Figure 8, we demonstrate sewing pattern editing guided by body pressure analysis, including adjustments to the cuff (a), upper bodice (b), lower bodice (c), and collar (d). As shown in the examples, our system accurately identifies areas with excessive tension and adjusts the corresponding sewing patterns to enhance comfort while preserving the overall design.
Generating New Garment Programs. A major challenge in traditional parametric pattern-making is the need to abstract symbolic programs for new sewing patterns, which demands both advanced programming skills and pattern-making expertise. Design2GarmentCode addresses this by correlating GarmentCode grammar with LMMs’ embedded pattern-making knowledge, enabling the automatic creation of new garment components. Figure 9 shows a layered-skirt component generated by our DSL-GA, along with 3D garment representations demonstrating different design parameters, such as skirt length (c), number of layers (d), and layer differences such as length difference and ruffling factor (e). The results demonstrate that our system consistently produces high-quality garment components that meet professional standards, while significantly reducing the time and expertise required to create new sewing pattern programs.
6 Conclusion
Design2GarmentCode transforms multi-modal design concepts into precise sewing patterns using LMMs to synthesize parametric programs. It addresses challenges related to data requirements, computation, and the limited precision of neural network-based methods. The experimental results demonstrate the system’s ability to capture design details while maintaining structural integrity and geometric precision in generated patterns.
Figure 10: Limitations of Design2GarmentCode, including failed to modeling thin structures like halter-neck, unable to model unconventional bodices and stitching relationships are limited to one-to-one mapping.
Despite these advantages, Design2GarmentCode currently cannot substantially alter GarmentCode’s underlying structure and logistics, which impacts generation quality due to inherent limitations in GarmentCode’s design and modeling capabilities. For example, the range of upper garment patterns is limited, making it difficult to model personalized segmentations (Figure 10 (b)). Additionally, for designs like halter necks or strapless tops (Figure 10 (a)), GarmentCode cannot model fine straps, leading to potential simulation failures. These constraints restrict the system’s ability to accurately represent certain complex or customized garment designs.